Can I get help in solving this problem?
a. For classes of 222 students, find the mean and standard deviation for the number born on the 4th of July. Ignore leap years.
b. For a class of 222 students, would two be an unusually high number who was born on the 4th of July?
a. The value of the mean is μ =
(Round to six decimal places as needed.)

Answer by Boreal(15235)

probability is (1/365)
so mean is np=222/365 or 0.608219
variance is np(1-p)=0.608*(364/365)=0.606
sd is sqrt(V)=0.78
probability two are born on that day is within 2 sd s of the mean
it is 365C2*(1/365)^2*(364/365)^220
=0.2727, Not at all unlikely.
Can use 6 decimal places for the others; question wasn't clear to which it applied.